# Natural Language Dataset Generation Framework for Visualizations Powered by Large Language Models

### v1

@@ 3.2.4. Question
In general, we conduct chart question answering to facilitate decision-making (Kim et al., 2020). Thus, the process involves analyzing charts through a question-answering, which ultimately leads to a conclusion and informs the decision-making. To generate questions, we employ a reverse thought process. This entails first identifying the decisions that can be derived from the charts (i.e., (#K) What higher-level decision can be made by analyzing this chart?), followed by formulating a possible conclusion that leads to such a decision (i.e., (#K) What is a possible conclusion that can be reached from this decision?). Finally, we determine what needs to be analyzed (i.e., (#K) What specific value can be retrieved to reach this conclusion? What are the mathematical operations to reach the conclusion?). We generate non-visual lookup and compositional questions using the provided values and mathematical operations. To transform these into visual questions, we identify the necessary visual attributes and incorporate them into the generated questions (i.e., (#K) What visual attributes are required to paraphrase this question?). Finally, we formulate an open-ended question designed to lead to the same conclusion obtained in the previous step.
@@ 3.3.2. Score-based Paraphrasing
Our paraphrasing technique is inspired by a linear interpolation in the latent space for image generation and manipulation as demonstrated in many system and application papers (Mozaffari et al., 2022; Abdrashitov et al., 2020; Aoki et al., 2022; Ko et al., 2022). It enables a smooth transition from one expression to another by focusing on creating controllable and meaningful variations of a single sentence. We employ a five-point Likert-scale and assign one of the axes to each. Here, we focus on altering only the sentence’s syntax, while maintaining its meaning. In detail, we provide LLMs with a sentence (i.e., Example Sentence) and an explanation about one of the defined axes (i.e., Axis) and its two directions (i.e., Direction-1, Direction-2). We assign a specific value on a Likert scale ranging from one to five, to paraphrase the sentence as if it were spoken by a person using a language with a certain degree indicated by the score. Below is the template prompt we used to paraphrase on a single axis:
1{Axis}
2
4
5Sentence: {Example Sentence}
6Score: {Score}
This technique can be extended to scenarios involving multiple scores and multiple axes (refer to an example result with two axes in Table 3). Here is the template prompt for paraphrasing across two language axes:
1{Axis-1}
2{Axis-2}
3
7
8Sentence: {Example Sentence}
9Score-A: {Score-A}, Score-B: {Score-B}
4. Vega-Lite Dataset
# of URLs / Vega-Lite specs
(a) URLs crawled 67,789
(b) URLs w/o duplicate 18,420
(c) URLs w/ license 7,408
(d) Specs after manual inspection 1,981
Quantity # of specs 4,318 9,897 6,680 52 709 1,981
Average # of keys in a spec 24 15 15 15 37 54
Simple (key $\leq$ 16) 0 6,164 6,354 41 186 73
Medium (key $\leq$ 24) 4,318 3,733 326 10 170 199
Complex (key $\leq$ 41) 0 0 0 1 179 733
Extra complex (key >>> 41) 0 0 0 0 174 976
Average depth of JSON 4.00 3.00 3.48 3.13 5.01 5.19
Average branching factor 1.22 1.44 1.18 1.17 1.41 1.38
Diversity Total # of unique keys 24 12 18 31 275 362
Composite views 0 0 0 0 136 746
Interaction (e.g., zoom, pan) 0 0 0 0 188 1,010
# of chart types 6 6 4 2 10 10
We have collected a new set of 1,981 real-world Vega-Lite specifications. It was used to test our LLM-based NL generation framework on diverse and complex charts. In this section, we present the details of our data collection process.
4.1. Dataset Construction
4.1.1. Search Queries.
4.1.2. Inclusion and Exclusion Criteria.
4.1.3. Post-processing.
4.2. Quantitative Analysis
4.2.1. Benchmarks.
4.2.2. Quality Metrics.
4.2.3. Complexity Levels.
We observe that the existing criteria used to establish the complexity levels of chart datasets are somewhat subjective and may not possess broad applicability (Luo et al., 2021a; Kantharaj et al., 2022; Lundgard and Satyanarayan, 2021). Instead, we suggest using the number of keys as a criterion for categorizing the complexity levels of charts, particularly in the context of Vega-Lite specifications. This is because, as explained above, the number of properties increases proportionately to the number of keys in a specification. To establish the standard number of keys, we refer to the Vega-Lite example gallery dataset (Satyanarayan et al., 2023b) and calculate the quartiles (Q1, Q2, Q3) based on the distribution of the number of keys. These quartiles, specifically 16, 24, and 41, are utilized as reference points to divide the specifications’ level of complexity. For instance, a specification with a total number of keys less than or equal to 16 is classified as ‘simple’ complexity. Likewise, a specification with a total number of keys greater than 16 and less than or equal to 24 is classified as ‘medium’ complexity (Figure 3).
4.2.4. Composite View, Interactivity, and Chart Type Distribution.
4.2.5. Results.
@@ 5. Experiments
Table 7. Quantitative comparison of benchmarks and LLM-generated utterances and questions. Two type of metrics were adopted, cross-distribution, which is to compare the two distributions to get the similarity and difference, and within-distribution, which is to compare the diversity within a single distribution. Each NL dataset has come from 4 sources, gold standard or benchmarks, LLM, LLM+P (paraphrased), LLM+P2 (paraphrased with 2 axes). The best metric from all sources are bold, while the best metric in ours (LLM, LLM+P, LLM+P2) are underlines.
@@
In this section, we introduce quantitative analysis of our generated NL datsets, lexical analysis on generated utterances, and types of low-level tasks in generated questions.
@@ 5.1. Experimental Setup
Our experiment aims to investigate the effectiveness of our framework in generating diverse NL datasets from Vega-Lite specifications, with a focus on accuracy and diversity. To achieve this, we apply tailored metrics to each NL dataset, taking into account their different characteristics. Specifically, we evaluate accuracy for L1/L2 captions, as they primarily serve the objective conveyance of chart information. We assess the diversity of utterances and questions, as it is important to reflect inclusive language usage among individuals with different background. The results are presented in Table 6 and Table 7 where each type of NL dataset is classified with capital English letter (A-G).
@@ 5.1.1. Benchmarks
For utterances and questions, we utilized crowd-sourced NL datasets gathered in prior studies (Srinivasan et al., 2021; Kim et al., 2020) (F-BM and G-BM). In case of utterance dataset, we only used the singleton case, so it was 804 sentences instead of 893. However, when it comes to captions, we encountered a challenge in finding suitable benchmarks for comparing with different caption levels. Previous research employed bitmap images of charts (Lundgard and Satyanarayan, 2021; Tang et al., 2023), whereas our approach leverages Vega-Lite specifications. This difference in data format prevented us from making an exact comparison.
@@ 5.1.2. Gold Standard Datasets
Given that benchmarks mostly focus on simple and medium level complexity with confined diversity, we decided to make a gold standard dataset to test the generalizable performance of our framework over diverse and complex charts. We referred to previous works (Kong et al., 2014; Kim et al., 2018) that have demonstrated how to create gold standard datasets. We selected 48 Vega-Lite specifications (Figure 3) by stratified sampling, taking into account their complexity level and whether they included interaction or composite views. Subsequently, three visualization experts (first three authors) collaborated to develop three guidelines for generating utterances and questions. These guidelines were crafted by referring to relevant suggestions and guidelines from prior research (Kim et al., 2021; Srinivasan et al., 2021; Wang et al., 2022; Kim et al., 2020). We began by creating sample utterances and questions for the same chart using the initial drafts, and jointly revised each guideline by reviewing the generated NL datasets. After making consensus about the final guidelines Appendix C, we divided the charts into thirds, with each person tasked with generating NL datasets for their assigned charts. This resulted in 48 utterances (comprising 16 commands, 16 queries, and 16 questions) and 80 questions (including 16 non-visual lookup, 16 non-visual compositional, 16 visual lookup, 16 visual compositional, and 16 open-ended questions) per each expert. After one expert created NL datasets for the assigned charts, the other two individuals conducted verification to find any issues or errors within these generated NL datasets. In cases where issues or errors were detected, all three experts convened to discuss and reach a consensus on how to address them. This collaborative effort resulted in the generation of 144 utterances with three different phrasings and 240 questions categorized into five types (D-Gold and E-Gold).
@@ 5.1.3. LLM-generated Datasets
To generate our datasets, we used an official API of GPT4444https://platform.openai.com/docs/models/gpt-4 with the gpt-4-0613 model. We set the temperature to 0.0, to solely observe the influence of our paraphrasing technique on diversity. We used different prompt for generating each dataset and paraphrasing the generated NL datasets (see Appendix A and Section 3.3). To be more specific, we generated all types of chart semantics, captions, utterances, and questions for the 48 sample charts, as well as all types of utterances for 30 charts from the benchmark. This resulted in a total of 432 chart semantics (A-LLM), 48 L1 captions (B-LLM), 48 L2 captions (C-LLM), 144 utterances (D-LLM), and 240 questions (E-LLM) for the 48 sampled charts, and 90 utterances for the 30 benchmark charts (F-LLM). Since the benchmark (Kim et al., 2020) did not include open-ended questions, we generated only four types of questions. This led to a total of 208 questions for the 52 charts (G-LLM).
We augmented our NL datasets for utterances and questions using the generated NL datasets (*-LLM) and the score-based paraphrasing technique, resulting in augmented paraphrased NL datasets (*-LLM+P and *-LLM+P2). With four language axes and five Likert-scale values (1-5), it is possible to generate 20 different versions (4*5) of paraphrased sentences for each original sentence (i.e., LLM+P). Likewise, in case of two axes, there are six combinations chosen from the four axes. Since there are five Likert-scale options for each axis, this leads to the generation of 150 (6*5*5) different paraphrased sentence versions per original sentence (i.e., LLM+P2). We meticulously generated all possible paraphrases and selected five distinct sets of NL datasets to mitigate any sampling bias. Thus we calculated metrics and their averages and standard deviations across these five sample sets.
When sampling the paraphrased sentences, our goal is to compare the syntactic diversity of different NL datasets while aligning the semantic diversity of the two datasets being compared to ensure a fair comparison. To this end, we adjust the frequency of each chart-NL pair in both datasets. This is necessary because the benchmark data exhibit biases in NL sentence distribution for each chart. For instance, one chart has 30 associated questions, while another chart has only one question. We count the frequency of each chart-NL pair and reflect the same frequency when augmenting the datasets. This became an issue when creating G-LLM+P, since one chart has 30 questions, which exceeds the maximum number of paraphrases possible (limited to 20) through our single-axis paraphrasing method. As a result, our overall number of NL datasets reaches 619.
Last, we included open-ended questions in E-LLM+P and E-LLM+P2, as these questions were available in E-Gold. However, we did not include them in G-LLM+P and G-LLM+P2 in Table 7 to preserve the semantic diversity of the datasets.
@@ 5.2. Quantitative Results
We next report the diversity of utterance and question. In terms of cross-distribution metrics, LLM+P exhibited the highest quality in terms of precision (D), precision and recall (F), and precision and FD (G). In case of datasets containing five question types (E), the metric results were not consistent. Specifically, LLM+P performed the best in FD, LLM was the best for precision, and LLM+P2 achieved the highest recall. When considering within-distribution metrics, LLM-generated and paraphrased datasets demonstrated greater diversity compared to the gold standard and benchmark datasets. On average, higher diversity was observed in 4.75 out of six metrics. For both question and utterance datasets (E, F), paraphrased datasets with two axes demonstrated greater diversity than paraphrased datasets with one axis in four out of six metrics. Conversely, in the other two datasets (D, G), paraphrased datasets with one axis exhibited higher diversity in four out of six metrics. In the utterance dataset (D), paraphrasing increased diversity in four out of six metrics, whereas in the question dataset (E), paraphrasing reduced diversity in four metrics. A summary of the results is presented in Table 7.
@@ 5.3. Lexical Analysis in Utterances
To gain a deeper understanding of the syntactic diversity in LLM-generated datasets, we conducted a lexical analysis on three NL datasets (F-BM, F-LLM+P, F-LLM+P2) to investigate the types of words used within each dataset. Our pre-processing steps encompassed sentence tokenization, converting all text to lowercase, removal of stopwords, and lemmatization. As evidenced by the quantitative outcome presented in the previous section, the LLM+P exhibited a notable richness in its lexical diversity. It contained a total of 555 unique words, surpassing the benchmark dataset’s count of 349 unique words. Also, the total word count in the LLM+P, amounting to 7,132 words, exceeded that of the benchmark dataset, which consisted of 4,480 words. In case of LLM+P2, it demonstrated an even greater number of unique words, totaling 608, surpassing both the benchmark and LLM+P datasets in this regard. However, the overall word count in LLM+P2 was lower at 6,645 words compared to the LLM+P dataset (7,132 words).

### v3

@@ 4.2.4. Question
In general, we conduct chart question answering to facilitate decision-making (Kim et al., 2020). Thus, the process involves analyzing charts through a question-answering, which ultimately leads to a conclusion and informs the decision-making. To generate questions, we employ a reverse thought process. This entails first identifying the decisions that can be derived from the charts (i.e., ($\mathcal{K}$) What higher-level decision can be made by analyzing this chart?), followed by formulating a possible conclusion that leads to such a decision (i.e., ($\mathcal{K}$) What is a possible conclusion that can be reached from this decision?). Finally, we determine what needs to be analyzed (i.e., ($\mathcal{K}$) What specific value can be retrieved to reach this conclusion? What are the mathematical operations to reach the conclusion?). We generate non-visual lookup and compositional questions using the provided values and mathematical operations. To transform these into visual questions, we identify the necessary visual attributes and incorporate them into the generated questions (i.e., ($\mathcal{K}$) What visual attributes are required to paraphrase this question?). Finally, we formulate an open-ended question designed to lead to the same conclusion obtained in the previous step.
@@ 4.3.2. Score-based Paraphrasing
Our paraphrasing technique is inspired by a linear interpolation in the latent space for image generation and manipulation as demonstrated in many system and application papers (Mozaffari et al., 2022; Abdrashitov et al., 2020; Aoki et al., 2022; Ko et al., 2022). This technique enables a smooth transition from one expression to another by focusing on creating controllable and meaningful syntactic variations of a single sentence. The key idea is that we assign language axesand employ a five-point Likert-scale to each. Here, we focus on altering only the sentence’s syntax, while maintaining its meaning. In detail, we provide LLMs with a sentence we want to paraphrase, and an explanation about one of the defined axes and its two directions. We assign a specific value on a Likert scale ranging from one to five, to paraphrase the sentence as if it were spoken by a person using a language with a certain degree indicated by the score. This technique can be extended to involve multiple axes and scores (refer to an example result with two axes in Table 5). The detailed prompts we used are presented in Appendix C.
@@ 5. Experiments
Table 7. Quantitative comparison of benchmarks and LLM-generated utterances and questions. Two type of metrics were adopted, cross-distribution, which is to compare the two distributions to get the similarity and difference, and within-distribution, which is to compare the diversity within a single distribution. Each NL dataset has come from 4 sources, gold standard or benchmarks, LLM, LLM.P (paraphrased), LLM.P2 (paraphrased with 2 axes). The best metric from all sources are bold, while the best metric in ours (LLM, LLM.P, LLM.P2) are underlines.
@@
In this section, we introduce quantitative analysis of our generated NL datasets, lexical analysis on generated utterances, and types of low-level tasks in generated questions.
@@ 5.1. Experimental Setup
Our experiment aims to investigate the effectiveness of our framework in generating diverse NL datasets from Vega-Lite specifications, with a focus on accuracy and diversity. To achieve this, we apply tailored metrics to each NL dataset, taking into account their different characteristics. L1/L2 captions are independent of the perception of humans or machines because they focus on conveying objective information (Lundgard and Satyanarayan, 2021). Thus, we measured accuracy to determine how precisely each caption level contained relevant information. We assess the diversity of utterances and questions, as it is important to reflect inclusive language usage among individuals with different background. The results are presented in Table 6 and Table 7 where each type of NL dataset is classified with capital English letter (A-G).
@@ 5.1.1. Benchmarks
For utterances and questions, we utilized crowd-sourced NL datasets gathered in prior studies (Srinivasan et al., 2021; Kim et al., 2020) (F-BM and G-BM). In case of utterance dataset, we only used the singleton case, so it was 804 sentences instead of 893. However, when it comes to captions, we could not find suitable benchmarks for comparing with different caption levels. Previous research employed bitmap images of charts (Lundgard and Satyanarayan, 2021; Tang et al., 2023), whereas our approach leverages Vega-Lite specifications. This difference in data format prevented us from making an exact comparison.
@@ 5.1.2. Gold Standard Datasets
Given that benchmarks mostly focus on simple and medium level complexity with confined diversity, we decided to make a gold standard dataset to test the generalizable performance of our framework over diverse and complex charts. We referred to previous works (Kong et al., 2014; Kim et al., 2018) that have demonstrated how to create gold standard datasets. We selected 48 Vega-Lite specifications (Figure 2) by stratified sampling, taking into account their complexity level and whether they included interaction or composite views. Subsequently, three visualization experts (first three authors) collaborated to develop three guidelines for generating utterances and questions. These guidelines were crafted by referring to relevant suggestions and guidelines from prior research (Kim et al., 2021; Srinivasan et al., 2021; Wang et al., 2022; Kim et al., 2020). We began by creating sample utterances and questions for the same chart using the initial drafts, and jointly revised each guideline by reviewing the generated NL datasets. After making consensus about the final guidelines Appendix D, we divided the charts into thirds, with each person tasked with generating NL datasets for their assigned charts. This resulted in 48 utterances (comprising 16 commands, 16 queries, and 16 questions) and 80 questions (including 16 non-visual lookup, 16 non-visual compositional, 16 visual lookup, 16 visual compositional, and 16 open-ended questions) per each expert. After one expert created NL datasets for the assigned charts, the other two individuals conducted verification to find any issues or errors within these generated NL datasets. In cases where issues or errors were detected, all three experts convened to discuss and reach a consensus on how to address them. This collaborative effort resulted in the generation of 144 utterances with three different phrasings and 240 questions categorized into five types (D-Gold and E-Gold).
@@ 5.1.3. LLM-generated Datasets
To generate our datasets, we used an official API of GPT4444https://platform.openai.com/docs/models/gpt-4 with the gpt-4-0613 model. We set the temperature to 0.0, to solely observe the influence of our paraphrasing technique on diversity. We used different prompt for generating each dataset and paraphrasing the generated NL datasets (see Appendix A and Appendix C). Here, we generated all types of chart semantics, captions, utterances, and questions for the 48 sample charts, as well as all types of utterances for 30 charts from the benchmark. This resulted in a total of 432 chart semantics (A-LLM), 48 L1 captions (B-LLM), 48 L2 captions (C-LLM), 144 utterances (D-LLM), and 240 questions (E-LLM) for the 48 sampled charts, and 90 utterances for the 30 benchmark charts (F-LLM). Since the benchmark (Kim et al., 2020) did not include open-ended questions, we generated only four types of questions. This led to a total of 208 questions for the 52 charts (G-LLM).
We augmented our NL datasets for utterances and questions using the generated NL datasets (*-LLM) and the score-based paraphrasing technique, resulting in augmented paraphrased NL datasets (*-LLM.P and *-LLM.P2). With four language axes and five Likert-scale values (1-5), it is possible to generate 20 different versions (4*5) of paraphrased sentences for each original sentence (i.e., LLM.P). Likewise, in case of two axes, there are six combinations chosen from the four axes. Since there are five Likert-scale options for each axis, this leads to the generation of 150 (6*5*5) different paraphrased sentence versions per original sentence (i.e., LLM.P2). We meticulously generated all possible paraphrases and selected five distinct sets of NL datasets to mitigate any sampling bias. Thus we calculated metrics and their averages and standard deviations across these five sample sets.
When sampling the paraphrased sentences, our goal is to compare the syntactic diversity of different NL datasets while aligning the semantic diversity of the two datasets being compared to ensure a fair comparison. To this end, we adjust the frequency of each chart-NL pair in both datasets. This is necessary because the benchmark data exhibit biases in NL sentence distribution for each chart. For instance, one chart has 30 associated questions, while another chart has only one question. We count the frequency of each chart-NL pair and reflect the same frequency when augmenting the datasets. This became an issue when creating G-LLM.P, since one chart has 30 questions, which exceeds the maximum number of paraphrases possible (limited to 20) through our single-axis paraphrasing method. As a result, our overall number of NL datasets reaches 619.
Last, we included open-ended questions in E-LLM.P and E-LLM.P2, as these questions were available in E-Gold. However, we did not include them in G-LLM.P and G-LLM.P2 in Table 7 to preserve the semantic diversity of the datasets.
@@ 5.2. Quantitative Results
We next report the diversity of utterance and question. In terms of cross-distribution metrics, LLM.P exhibited the highest quality in terms of precision (D), precision and recall (F), and precision and FD (G). In case of datasets containing five question types (E), the metric results were not consistent. Specifically, LLM.P performed the best in FD, LLM was the best for precision, and LLM.P2 achieved the highest recall. When considering within-distribution metrics, LLM-generated and paraphrased datasets demonstrated greater diversity compared to the gold standard and benchmark datasets. On average, higher diversity was observed in 4.75 out of six metrics. For both question and utterance datasets (E, F), paraphrased datasets with two axes demonstrated greater diversity than paraphrased datasets with one axis in four out of six metrics. Conversely, in the other two datasets (D, G), paraphrased datasets with one axis exhibited higher diversity in four out of six metrics. In the utterance dataset (D), paraphrasing increased diversity in four out of six metrics, whereas in the question dataset (E), paraphrasing reduced diversity in four metrics. A summary of the results is presented in Table 7.
@@ 5.3. Lexical Analysis in Utterances
To gain a deeper understanding of the syntactic diversity in LLM-generated datasets, we conducted a lexical analysis on three NL datasets (F-BM, F-LLM.P, F-LLM.P2) to investigate the types of words used within each dataset. Our pre-processing steps encompassed sentence tokenization, converting all text to lowercase, removal of stopwords, and lemmatization. As evidenced by the quantitative outcome presented in the previous section, the LLM.P exhibited a notable richness in its lexical diversity. It contained a total of 555 unique words, surpassing the benchmark dataset’s count of 349 unique words. Also, the total word count in the LLM.P, amounting to 7,132 words, exceeded that of the benchmark dataset, which consisted of 4,480 words. In case of LLM.P2, it demonstrated an even greater number of unique words, totaling 608, surpassing both the benchmark and LLM.P datasets in this regard. However, the overall word count in LLM.P2 was lower at 6,645 words compared to the LLM.P dataset (7,132 words).
@@ 5.4. Types of Low-level Tasks in Questions
6. Application
6.1. Finetuning LLMs for Data Visualization
Table 9. The result of finetuning experiment. The LLMs trained with NL datasets generated by our framework either matched or surpassed the performance of LLMs (C, D, E) compared to when using only the benchmark dataset (A).
Source Train # Test # Accuracy (#)
A. BM (Srinivasan et al., 2021) 723 81 76.3% (61.8)
B. LLM.P 723 81 58.8% (47.6)
C. BM + LLM.P 723 81 76.8% (62.2)
D. BM + LLM.P 1446 81 83.2% (67.4)
E. BM + LLM.P + LLM.P2 2169 81 85.4% (69.2)
We demonstrate that the NL datasets generated by our framework can be used to augment the performance of ML models. It is important to note, however, that the effectiveness of our datasets is contingent upon the availability of a sufficient number of human-generated datasets that exhibit similar distributions to the test datasets. We believe our approach serves as a cost-effective and efficient way to be used in conjunction with the conventional method of crowdsourcing human-generated NL datasets. In essence, our framework’s output can be strategically employed as supplementary datasets for finetuning LLMs.
To be specific, to replicate the benchmark dataset’s experiment (Srinivasan et al., 2021), we performed an experiment to classify ten chart types (e.g., colored scatterplots, stacked & grouped bar charts, multiseries line charts, etc.) using utterances. This classification task is important as it can be further used for building visualization systems like chart type recommendation. We prepared five datasets for finetuning: A. the benchmark dataset (723 utterances), B. the utterances generated and paraphrased with one axis by our framework (723 utterances), C. half of A and half of B (362 from A + 361 from B), D. A and B (723 from A + 723 from B), E. D as well as the utterances generated and paraphrased with two axes by our framework (1446 from D + 723 additional utterances). Only 90% of the benchmark dataset is used for finetuning and the rest 10% were used for the test. Similarly, we used only 90% of our datasets to maintain an equal number of utterances as in the benchmark dataset. Following common ML practices, we selected OpenAI’s babbage-002 model for training smaller models on downstream tasks, setting hyperparameters to default configurations (i.e., number of epochs as 3, learning rate multiplier as 2, and the batch size were 1, 1, 1, 2, 4 for each case). Each experiment was repeated five times to calculate the average accuracy to mitigate the stochastic behavior of LLMs.
As denoted in Table 9, we observed an increase in performance when using LLM-generated NL datasets alongside the benchmark dataset for finetuning the models. Using only the benchmark datasets resulted in an accuracy of 76.3% (61.8 accurate prediction on average out of 81, Table 9-A). When we combined the benchmark datasets with our dataset, the accuracy slightly improved to 76.8% (62.2 out of 81, see Table 9-C), indicating that the addition of a non-human-generated datasets did not negatively impact accuracy. Moreover, the performance increased to 83.2% when we leveraged additional NL datasets generated by LLMs (67.4 out of 81, Table 9-D). The accuracy was the highest when we used more NL datasets paraphrased with two language axes by our framework, which is 85.4% (69.2 out of 81, Table 9-E). Last, using only LLM-generated NL datasets showed decreased accuracy, which is 58.8% (47.6 out of 81, Table 9-B).
The results suggest that using the NL datasets, generated and paraphrased by our framework, can enhance the performance of ML models in downstream tasks. We believe a key factor in this improved performance is the increased syntactic diversity of the generated utterances, which also accurately mimic semantic characteristics. Our results align with a previous finding that utilizing AI-generated datasets can become a more cost-effective strategy for training scalable ML models with significantly fewer human labels (Bai et al., 2022). This suggests that the synergistic use of both human efforts and our automated framework can substantially enhance the quality of training data and the performance of the models.
6.2. Leveraging Fully-automatic and Mixed-initiative Modes in VL2NL
Figure 4. A system with two modes (fully-automatic and mixed-initiative) to generate NL datasets using VL2NL. The mixed-initiative mode encompasses several features. First, users can select the types of NL datasets they want to generate (C). They can inspect each chart (A) and subsequently choose the specific ones they wish to use for generating NL datasets (E). Users can change or provide information that the system utilizes (B). Once these are completed, the system returns the generated NL datasets (D). In contrast, the fully-automatic mode does not include (B). As a result, dataset generation in this mode strictly follows the scaffolding defined by researchers, along with key questions and answers generated by LLMs.
To further explore how visualization researchers can use our framework, we performed a case study with two experts: E1, a professor, and E2, a postdoctoral researcher. Both have earned their Ph.D. in visualization and have conducted research for 10 and 7 years, respectively.
For the case study, we implemented a system with two modes: fully-automatic and mixed-initiative (Figure 4). In the fully-automatic mode, the scaffolding is set by us and key questions are generated automatically by the LLMs. Therefore, users had no control, but could click the button to generate NL datasets for their chosen charts. In the mixed-initiative mode, users can select which scaffolding to consider and provide additional information as answers to key questions. They can actively contribute by specifying directions to steer its focus accordingly. For example, in case of L1 caption, they can choose which chart semantics to consider or add more when generating it. Similarly, for question, users can make a high-level decision themselves, specifying where or what to focus on when analyzing the charts.
To clarify our study protocol, we first provided the experts with an overview of our framework’s concept. Following this, they were given a task to create NL utterances for 10 line charts, which depicted stock prices of various technology companies (Xu and Cohen, 2018). This was conducted using two modes: fully-automatic and mixed-initiative. We emphasized to the participants that the utterances they generated would be instrumental in training an ML model to translate these utterances back into the corresponding line charts. We also highlighted the significance of utterance diversity in enhancing the performance of ML models, based on our discussion in (Section 6.1). Finally, the experts provided feedback on their anticipated use of both modes for generating utterances. Each expert spent approximately 45 minutes for the study.
Both experts agreed on using both modes to generate utterances more effectively for training ML models. Specifically, E1 suggested the following scenario: initially, researchers generate a large number of utterances automatically to observe their distribution. Next, they identify areas lacking in diversity, which then become the focus for generating additional utterances subsequently. By repeatedly testing and generating utterances in these sparse areas, particularly using the mixed-initiative mode, they can achieve a more diverse and evenly distributed utterances. This process, iterated over multiple times, could improve the performance of the ML models. Similarly, E2 also advocated starting with the fully-automatic mode before using the mixed-initiative mode. E2 said this approach allows experts to better understand the model’s behavior and the nature of the utterances it generates. This step is crucial to avoid ‘option paralysis,’ a state of cognitive overload that may occur when faced with a lot of choices without a clear strategy for improvement. With a deeper understanding of the model’s behavior, they can proceed more effectively.